A 100g bullet is fired on a ballistic gel with a force constant of 200N/m/s. The ballistic gel is large enough to cover the trajectory of the bullet. If the initial speed of the bullet is 700m/s upon entering the ballistic gel, how long (in seconds) will it take the gel to slow the bullet to 0.01m/s? Ignore acceleration due to gravity. 


1
Expert's answer
2021-03-31T07:16:13-0400

Solution.

"m=100g=0.1kg;"

"k=200N\/(m\/s);"

"v_o=700m\/s;"

"v=0.01m\/s;"

"t-?;"

"a=\dfrac{v-v_o}{t} \implies t=\dfrac{v-v_o}{a};"

"a=\dfrac{F}{m};" "k=\dfrac{F}{\Delta v}\implies F=k\Delta v;"


"t=\dfrac{\Delta v m}{k\Delta v}=\dfrac{m}{k};"


"t=\dfrac{0.1kg}{200N/(m/s)}=0.0005s;"


Answer: "t=0.0005s."